# Identifying compact symmetric objects with high-precision VLBI and Gaia astrometry

### v1

@@ 3 Results
Our application of the Gaia+VLBI method yielded definitive classifications for 32 of 40 CSO candidates, demonstrating the effectiveness of combining high–precision optical astrometry with radio imaging. We confirmed 22 sources as genuine CSOs (Fig 4), identified 10 as core–jet AGN (Fig 5), while 8 sources (Fig. 6) remain ambiguous due to significant radio–optical positional discrepancies.
@@ 3.1 Confirmed CSOs
Among 22 confirmed CSOs, 17 show direct Gaia+VLBI positional alignment, with 5 supported by additional morphological and spectral evidence despite larger positional offsets (Fig 4). These sources display diverse morphologies that provide crucial insights into their physical nature and evolutionary stages.
The majority exhibit classic double–lobed structures with symmetric hotspots, exemplified by J1111+1955 and J1326+3154. Several sources show complex morphologies that reveal rich jet–ISM interactions. For example, J0832+1832 and J1815+6127 feature multiple hotspots in their radio lobes and large misalignment, suggesting jet deflection or multiple impact sites at the jet–ISM interface. The dramatically bent jet in J1158+2450 presents compelling evidence of environmental influence, while the elongated, clumpy structure of J1244+4048 indicates jet instabilities driven by ISM entrainment.
Our multi-frequency approach proves crucial for reliable CSO identification. The 2.3-GHz observations of J0741+2706, J0832+1832, J1256+5652, and J2355+4950 reveal extended emission structures invisible at higher frequencies, providing critical evidence for their CSO classification. This frequency-dependent effect occurs because steep–spectrum components, particularly counter–lobes, become too faint to detect at 5 and 8.4 GHz due to the combined effects of spectral steepening and Doppler deboosting.
The spectral properties of our sample offer valuable insights into CSO physics. Most sources display the characteristic steep-spectrum lobes () consistent with aging relativistic electron populations. Central regions generally show flatter spectra, steepening toward the lobes, indicating continuous particle acceleration near the core. Of particular interest are the jet–ISM interaction regions, which often exhibit distinctly flat or even inverted spectra, as seen in the southern lobe of J0943+1702 and the northeastern jet of J1158+2450. These spectral features provide direct evidence of ongoing particle acceleration in interaction zones.
@@
Kinematic analysis (Table 2) reveals diverse hotspot separation speeds in our CSO sample. Most sources display sub-relativistic motions (), characteristic of symmetric jet expansion in young radio galaxies (Peck & Taylor, 2000; Kawakatu & Kino, 2006; An & Baan, 2012). However, two quasars (J0943+1702 and J1602+2418) show mildly relativistic hotspot advance speeds of and , respectively, though with substantial uncertainties. While J1111+1955, a galaxy-type CSO, appears to show an unusually high apparent speed, this measurement is based on limited data spanning only three epochs over three years and requires confirmation with longer-term monitoring. The measured proper motions yield kinematic age estimates spanning from years (J1111+1955) to over 1600 years (J1234+4753), revealing diverse evolutionary stages. Three sources (J1110+4817, J1256+5652, J1815+6127) show apparent negative speeds, likely arising from complex changes in emission structure or projection effects (An et al., 2012).
@@
Notably, five nearby () CSOs — J0048+3157 (NGC 262, Mrk 348), J1247+6723, J1407+2827, J1511+0518, and J1945+7055 — show systematic radio–optical offsets despite strong morphological and spectral evidence supporting their CSO nature. This includes well-studied sources like J1407+2827 (OQ 208, Stanghellini et al. 1997; An et al. 2012; Wu et al. 2013), where the Gaia position coincides with a radio lobe rather than the core, likely due to bright optical emission from jet–ISM interaction regions. These cases highlight how environmental effects in host galaxies can complicate optical position interpretation. In starburst or merging host galaxies, significant dust extinction and complex optical structures can lead to substantial Gaia positioning uncertainties. While illustrating limitations of the Gaia+VLBI method for nearby galaxies with weak or obscured optical nuclei, their CSO classification remains secure based on comprehensive morphological and spectral evidence.
Our confirmed CSO sample spans a broad range of redshifts (0.015 to 1.791, median 0.269) and projected linear sizes (3.6 to 290 pc), all below the conventional 1 kpc limit for CSOs (Table 1). This diversity facilitates investigation of different aspects of CSO evolution, from initial jet propagation to the development of large–scale radio structure.
@@ 3.2 Core–jet sources
We identified 10 sources as core–jet AGN based on distinctive characteristics that clearly differentiate them from CSOs (Fig. 5). The key identifier is the alignment of Gaia optical positions with the brightest components at the ends of the elongated radio structures, a hallmark of relativistically beamed emission.
A key phenomenon observed in these sources is the core shift effect — the systematic change in apparent core position with observing frequency due to synchrotron self-absorption gradients along the jet (Lobanov & Zensus, 1999). At higher frequencies, we observe regions closer to the SMBH, while optical emission originates even closer (Kovalev et al., 2008). In several core–jet sources (e.g., J0753+4231, J1311+1417), the optical peak associates with a weaker upstream component rather than the brightest radio feature, indicating the true core location obscured by synchrotron self-absorption at radio frequencies. In J0003+2129, the Gaia position coincides with the counter–jet region, which remains invisible due to Doppler deboosting. In most sources, the optical–radio offsets are less than 1 mas, consistent with astrometric uncertainties. An exception is J2022+6136, where the flat-spectrum radio component is offset by more than 13 mas from the Gaia position, likely associated with an optically bright hotspot downstream of the jet.
@@ 3.3 CSO candidates
Eight sources (Fig. 6) defy straightforward classification, as these objects exhibit significant discrepancies between their Gaia optical positions and VLBI radio structures, with positional offsets exceeding combined astrometric uncertainties. This suggests that Gaia detections may correspond to physically unrelated foreground/background objects or extended host galaxy features rather than the AGN cores (Bailer-Jones et al., 2019).
J0831+4608 and J1254+1856 have been observed only at a single epoch and frequency, limiting definitive classification. While their morphology suggests CSO characteristics, confirmation requires multi-epoch multi-frequency VLBI observations to reveal their spectral properties and proper motions.
J1559+5924 displays a triple structure with a central bright component and two extended features, resembling a mini FR I galaxy. It remains a CSO candidate, potentially belonging to a subclass of low-power CSOs with weak, two-sided jets. Further investigation of this source could provide insights into the diversity of CSO morphologies.
J0119+3210, J0906+4636, and J1148+5924 exhibit complex radio emission without the compact, edge-brightened lobes typical of CSOs. Their extended emission may result from nuclear star formation, supernova remnants, or diffuse jets. The significant offsets between their optical and radio emission centers suggest complex interactions between the AGN and host galaxy environment.
J0650+6001 challenges conventional classification schemes with its unusual combination of properties. While its Gaia position near the southernmost component suggests a core–jet source, it displays an overall gigahertz-peaked spectrum (GPS) and relatively slow jet motion (, Lister et al. 2019). However, its notable variability contradicts typical CSO characteristics, making its true nature uncertain.
These ambiguous cases manifest the limitations of current classification schemes and highlight the importance of comprehensive multi-criteria analysis in understanding the full diversity of compact radio sources. Their study may reveal new subclasses or evolutionary stages bridging established categories, advancing our understanding of radio galaxy evolution.
@@ 4.1 Efficacy and validation of the Gaia+VLBI method
The integration of Gaia optical astrometry with VLBI radio imaging provides a powerful new approach for identifying AGN cores in compact radio sources. Our successful confirmation of 22 CSOs demonstrates this method’s effectiveness in addressing a fundamental challenge in CSO identification. The key strengths of this method include precise core localization even when radio emission is weak or obscured, unambiguous discrimination between CSOs and core–jet sources, independent verification of radio structure orientation, and enhanced accuracy in kinematic measurements.
Our method complements recent comprehensive studies such as the ‘bona fide’ CSO catalogue (Kiehlmann et al., 2024a). While their catalogue of 79 CSOs, including 43 edge-brightened CSOs (their ‘CSO 2’ class) with redshifts, relies on strict radio-based criteria, our method introduces precise optical–radio correlation as an additional diagnostic tool. The 60% overlap (13 out of 22) between our confirmed CSOs and Kiehlmann’s catalogue validates our approach while highlighting potentials for new insights. Two sources we classify as core–jet (J1816+3457, J2022+6136) are listed as CSOs by Kiehlmann et al. (2024a), demonstrating the complexity of CSO identification.
However, several important limitations must be considered:
@@
Optical emission from jets or hotspots may dominate over core emission in some cases (e.g. J1407+2827), potentially leading to misidentification.
Reduced effectiveness for nearby galaxies where extended optical structure affects Gaia measurements (e.g. J1511+0518).
Potential bias against heavily obscured AGN due to systematic offsets from dust obscuration.
@@
To mitigate these limitations, we recommend a multi-wavelength approach combining data across different bands, careful assessment of host galaxy properties, and integration with traditional radio–based classification methods. Despite these constraints, the Gaia+VLBI method represents a significant advance in CSO identification, particularly in their earliest stages where traditional methods often prove inadequate. Future work will expand this pilot study by cross-matching Gaia with over 20,000 Astrogeo VLBI sources. This extensive dataset will illuminate selection effects, environmental factors, and CSO evolutionary pathways across diverse populations.
4.2 Radiation characteristics of CSOs
4.2.1 Radio Morphology and Spectral Properties
@@
The majority () of confirmed CSOs exhibit classical double–lobed structure with edge–brightened hotspots (e.g., J1111+1955, J1358+4737, J1602+2418, and J1823+7938) supporting theoretical models where young jets propagate symmetrically through dense interstellar medium. The emission typically becomes detectable when the jet terminates in a shock, efficiently converting kinetic energy into relativistic particles (Blandford & Rees, 1974; Begelman et al., 1984; Scheuer, 1995; Meisenheimer et al., 1989). For high-power jets, the kinetic energy is rarely lost as radiation during jet propagation, which often makes the core and inner jet invisible.
@@
Only six CSOs (J0741+2706, J1158+2450, J1234+4754, J1244+4048, J1256+5652, and J1310+3403) are detected with central radio components, which likely represent unresolved inner jet regions rather than cores. Several CSOs display complex morphologies indicating strong jet–ISM interactions. For example, J0832+1832 and J1815+6127 show multiple hotspots suggesting jet deflection or impact with dense ISM clouds (Mendoza & Longair, 2002), or jet head impacting multiple sites of the interface (Carvalho & O’Dea, 2002; An et al., 2012).
@@
The dramatically bent jet in J1158+2450 provides direct evidence of how strong jet–ISM interactions can sculpt the CSO morphology (Mukherjee et al., 2016). Similarly, the elongated, clumpy jet structure in J1244+4048 likely results from jet instabilities induced by entrainment of the ISM (e.g. Bicknell, 1984). In J1256+5652 (Mrk 231), we observe an east–west oriented jet within mas of the core, but extended emission at mas scales (Wang et al., 2021). This structure hints at multiple episodes of AGN activity, offering a rare glimpse into the duty cycle of these young radio sources and challenging simple, monotonic evolution models (Saikia & Jamrozy, 2009; Brocksopp et al., 2011; Orienti & Dallacasa, 2020; Brocksopp et al., 2007). Such recurrent activity (Section 4.2.3), observed in numerous compact radio sources, reveals key aspects of both AGN feedback processes and long-term evolution of radio galaxies (Schoenmakers et al., 2000; Konar et al., 2012; Nandi et al., 2019).
@@
The spectral properties provide additional information on the physical conditions in these sources. The overall radio spectra typically show a convex shape with a turnover near GHz, attributed to synchrotron self-absorption in compact, high-density regions (O’Dea, 1998). This characteristic is not limited to our low-redshift sample but also observed in high-redshift CSOs (Sotnikova et al., 2021).
The central regions generally exhibit flatter spectra, steepening towards the lobes (e.g., J1234+4753 and J1244+4048), consistent with theoretical models of continuous particle acceleration in central regions, followed by spectral aging as the plasma flows outward (Pacholczyk, 1970). Regions of jet–ISM interaction often show flat or inverted spectra (e.g., the southern lobe of J0943+1702), indicating ongoing particle acceleration in these interaction zones (Godfrey et al., 2009).
@@ 4.2.2 Brightness temperatures and Doppler boosting
Our VLBI observations reveal a distinct contrast between CSOs and other radio-loud AGN classes in their relativistic beaming properties. CSO components typically exhibit moderate brightness temperatures ( K), well below the inverse Compton limit of K (Kellermann & Pauliny-Toth, 1969; Readhead, 1994). This, combined with derived Doppler-boosting factor , distinguishes CSOs from highly beamed sources like blazars (Kovalev et al., 2005; Jorstad et al., 2005; Homan et al., 2006; Hovatta et al., 2009; Liodakis et al., 2018; Cheng et al., 2020). Unlike flat-spectrum radio quasars with core K and Doppler factors (Pushkarev & Kovalev, 2012), our CSO sample consistently fall below these thresholds (Readhead et al., 1996b; Polatidis & Conway, 2003; An & Baan, 2012), reinforcing their interpretation as young, intrinsically symmetric sources.
The combination of moderate brightness temperatures and minimal Doppler boosting in CSOs allow direct measurement of intrinsic source properties without the complicating effects of relativistic beaming, making CSOs as critical probes of fundamental AGN jet physics and early evolution.

### v2

@@ 3 Results
Our application of the Gaia+VLBI method yielded classifications for 30 of 40 CSO candidates, demonstrating the effectiveness of combining high–precision optical astrometry with radio imaging. We confirmed 20 sources as bona fide CSOs, and identified ten as core–jet AGN (Figures S2–S3, Tables S3–4). The remaining ten sources (Fig. S4, Table S5) are ambiguous due to significant radio–optical positional discrepancies.
@@ 3.1 Confirmed CSOs
Among the 20 confirmed CSOs, 15 show direct Gaia+VLBI alignment; the remaining five are well studied in the literature and are confirmed as CSOs by independent evidence, despite larger Gaia–VLBI offsets.. Most exhibit classic double-lobed structures (e.g., J1111+1955, J1326+3154), while several show complex morphologies indicating jet-ISM interactions. Sources such as J0832+1832, J1158+2450, and J1815+6127 display asymmetric brightness where components closest to the core appear brightest, suggesting strong environmental influence. Additional evidence of the environment includes jet deflection (J0832+1832, J1815+6127), bending (J1158+2450), and clumpy morphology (J1244+4048). Multifrequency coverage is essential: 2.3-GHz data reveal extended, steep-spectrum features in J0741+2706, J0832+1832, J1256+5652, and J2355+4950 that are too faint at higher frequencies owing to spectral steepening and Doppler deboosting.
Lobes generally have steep spectra () from aging relativistic electrons. Central regions generally show flatter spectra, steepening toward the lobes, consistent with opacity and/or ongoing particle injection near the core, followed by spectral aging as the plasma flows outward. Hotspots in young radio sources are expected to be optically thin with steep spectra (). Flat or inverted spectral indices in some regions (e.g., NE jet of J1158+2450) are best explained by opacity effects (synchrotron or free-free absorption) or contamination from unresolved core/inner jet emission (Taylor et al., 1996b; Orienti & Dallacasa, 2014).
@@
Kinematic analysis (Table 2) reveals diverse hotspot separation speeds in our CSO sample. Most sources display sub-relativistic motions (), which are characteristic of symmetric jet expansion in young radio galaxies. J1602+2418 shows mildly relativistic speed of , though with substantial uncertainties. J1111+1955 may exhibit an unusually high apparent speed that requires confirmation with longer time baseline. Estimated kinematic ages span years (J1602+2418) to over 1600 years (J1234+4753), revealing diverse evolutionary stages. Three sources (J1110+4817, J1256+5652, J1815+6127) show apparent negative speeds, likely from structure changes or projection effects (An et al., 2012).
@@
Five nearby () CSOs — J0048+3157 (NGC 262, Mrk 348), J1247+6723, J1407+2827, J1511+0518, and J1945+7055 — show large radio–optical offsets despite characteristic CSO morphology and spectra. In J1407+2827 (OQ 208, Stanghellini et al. 1997; An et al. 2012; Wu et al. 2013), the Gaia position coincides with a radio lobe, likely due to bright optical emission from a jet–ISM interaction site. Host-galaxy factors (central asymmetries, host spectral components, and dust extinction) can shift optical centroids relative to VLBI positions (Plavin et al., 2019), and starburst/merging hosts can further increase Gaia uncertainties. These cases illustrate limitations of the Gaia+VLBI method for nearby, optically weak or obscured nuclei, but their CSO classification remains secure from morphology and spectra.
The projected linear sizes of the confirmed CSOs range from 3.6 to 290 pc, all within the canonical kpc CSO scale (Table 4). This diverse distribution enables tests of CSO evolution from initial jet launch to early growth of radio structure.
@@ 3.2 Core–jet sources
We identified ten core–jet sources in which Gaia positions coincide with the brightest terminal component of an elongated radio structure, which is consistent with relativistic beaming (Fig. S3). A frequency-dependent core shift (systematic position changes with observing frequency due to synchrotron self-absorption gradients along the jet) modulates these alignments (Lobanov & Zensus, 1999). Higher radio frequencies probe regions closer to the SMBH, while the optical centroid typically lies even farther upstream (Kovalev et al., 2008). In several sources (e.g., J0753+4231, J1311+1417), the optical position aligns with a weaker upstream component rather than the brightest radio feature, marking the true core locations. In J0003+2129, the Gaia position is displaced toward the counter-jet relative to the radio peak; the absence of radio emission at the optical position is consistent with strong opacity at the jet base, whereas the optical traces the unobscured nucleus. Most optical-radio offsets are mas and compatible with astrometric uncertainties. The mas offset for J2022+6136 likely reflects its faintness and large astrometric uncertainty (). Kinematic analysis (Table S6) shows apparent jet speeds of –, well above the subrelativistic CSO hotspot motions, confirming their relativistic jets.
@@ 3.3 CSO candidates
Ten sources show Gaia–VLBI offsets larger than the combined astrometric uncertainties, implying that the Gaia detections may trace unrelated foreground/background objects or extended host features rather than the AGN cores (Bailer-Jones et al., 2019). J0831+4608 and J1254+1856 were observed at only one epoch and frequency; their morphologies are CSO-like, but confirmation requires multi-epoch, multifrequency VLBI to measure spectra and proper motions. J0943+1702 shows a symmetric double at 2.3 GHz with the Gaia position near the midpoint. At 8.7 GHz the structure resolves into N–C–S, with C having flat-spectrum and N/S steep-spectrum. Total-flux variability appears modest (S band), but the apparent hotspot speed is atypical of CSOs and may reflect projection or pattern speeds; we therefore retain a conservative CSO-candidate label. J1823+7938 displays a double-lobe morphology at 2.3 GHz; at 8.4 GHz, the Gaia centroid lies near the E component, which shows flat spectrum, favouring a core+jet geometry. Its roughly symmetric morphology, slow (subluminal) expansion, and low S-band variability are all consistent with a CSO, but the Gaia offset from the geometric midpoint motivates a conservative CSO-candidate classification.
J1559+5924 shows a bright central component with two diffuse extensions, resembling a mini FR I; it remains a CSO candidate, possibly a low-power, two-sided-jet system. J0119+3210, J0906+4636, and J1148+5924 display complex emission without compact, edge-brightened lobes. Their component brightness temperatures (– K) and radio powers (– W Hz-1) exceed thermal star-formation values yet fall below typical blazar cores, consistent with non-thermal, low-power jets rather than classical CSOs. Also J0650+6001 remains ambiguous: its Gaia position near the southern component suggests a core–jet source, yet the overall GHz-peaked spectrum and slow jet motion () are CSO-like (Orienti & Dallacasa, 2010; An et al., 2012; Lister et al., 2019); its strong total-flux variability (from 1000 mJy in 2012 to 200 mJy in 2021) argues against a standard CSO. These cases highlight classification limitations and the need for a multi-criterion approach. Targeted, multi-epoch and multiband VLBI will be decisive for confirming or rejecting CSO status.
@@ 4.1 Efficacy and validation of the Gaia+VLBI method
Combining Gaia astrometry with VLBI imaging enables robust core identification in compact radio sources. Our confirmation of 20 CSOs demonstrates this method’s effectiveness, providing core localization even when radio cores are faint or obscured, clear separation of CSOs from core–jet systems, independent checks on structure orientation, and improved kinematic baselines.
This method complements radio-only catalogues such as the recent bona fide CSO sample (Kiehlmann et al., 2024a, b): while that catalogue of 79 CSOs, including 43 edge-brightened CSOs (CSO 2 class) with redshifts, applies strict radio-based criteria, our optical–radio registration provides an additional diagnostic. The overlap (12/19) between our confirmed CSOs and their list supports the approach. Two cases that are not consistent with the other classifications are classified as core-jet sources (J1816+3457 and J2022+6136), which highlights the known ambiguities in CSO identification. Several important limitations must be considered:
@@
Effectiveness decreases for nearby galaxies with complex optical structure.
Bias against heavily dust-obscured or optically faint nuclei.
@@
To mitigate these limitations, we recommend a multiwavelength approach combining data across different bands, careful assessment of host galaxy properties, and integration with traditional radio–based classification methods. Despite caveats, Gaia+VLBI advances early-stage CSO identification where radio-only methods can be inconclusive. Future work will scale to a -source cross-match with Astrogeo VLBI, enabling quantitative tests of selection effects, environments, and CSO evolution across diverse populations.
4.2 Radio morphology and spectral properties
@@
The majority () of confirmed CSOs exhibit a classical double-lobed structure with edge-brightened hotspots (e.g., J1111+1955, J1358+4737, J1602+2418), supporting theoretical models where young jets propagate symmetrically through a dense ISM. The emission typically becomes detectable when the jet terminates in a shock, efficiently converting kinetic energy into relativistic particles (Blandford & Rees, 1974; Begelman et al., 1984; Scheuer, 1995; Meisenheimer et al., 1989). For high-power ( W Hz-1) jets, kinetic energy is rarely lost as radiation during jet propagation, which often makes the core and inner jet invisible.
@@
Only six CSOs (J0741+2706, J1158+2450, J1234+4753, J1244+4048, J1256+5652, and J1310+3403) are detected with central radio components, which likely represent unresolved inner jet regions rather than cores. Several CSOs display complex morphologies indicating strong jet–ISM interactions. For example, J0832+1832 and J1815+6127 show multiple hotspots suggesting jet deflection or impact with dense ISM clouds (Mendoza & Longair, 2002), or jet head impacting multiple sites at the interface (Carvalho & O’Dea, 2002; An et al., 2012).
@@
The dramatically bent jet in J1158+2450 provides direct evidence of how strong jet–ISM interactions can sculpt the CSO morphology (Mukherjee et al., 2016). Similarly, the elongated, clumpy jet structure in J1244+4048 likely results from jet instabilities induced by entrainment of the ISM (e.g., Bicknell, 1984). In J1256+5652 (Mrk 231), we observe an east-west oriented jet within mas of the core, but extended emission at mas scales (Wang et al., 2021), hinting at multiple episodes of AGN activity and challenging simple, monotonic evolution models (Saikia & Jamrozy, 2009; Brocksopp et al., 2011; Orienti & Dallacasa, 2020; Brocksopp et al., 2007). Such recurrent activity (Section 4.2.2), observed in numerous compact radio sources, reveals key aspects of both AGN feedback processes and long-term evolution of radio galaxies (Schoenmakers et al., 2000; Konar et al., 2012; Nandi et al., 2019).
@@
The spectral properties provide additional information on the physical conditions in these sources. The overall radio spectra typically show a convex shape with a turnover near GHz, attributed to synchrotron self-absorption in compact, high-density regions (O’Dea, 1998). This characteristic is observed in both low-redshift () and high-redshift () CSOs (Sotnikova et al., 2021).
The central regions generally exhibit flatter spectra, steepening toward the lobes (e.g., J1234+4753 and J1244+4048); this is consistent with theoretical models of continuous particle acceleration in central regions and is followed by spectral aging as the plasma flows outward (Pacholczyk, 1970). The observed morphological and spectral diversity indicates that CSOs are not merely scaled-down versions of larger radio galaxies (An & Baan, 2012), they represent crucial laboratories for studying jet-ISM interactions, shock physics, and AGN feedback in action (Tadhunter, 2016).
@@ 4.2.1 Brightness temperatures and Doppler boosting
Our VLBI observations reveal a distinct contrast between CSOs and other radio-loud AGN classes in their relativistic beaming properties. CSO components typically exhibit moderate brightness temperatures ( K) well below the inverse Compton limit of K (Kellermann & Pauliny-Toth, 1969; Readhead, 1994). This, combined with the derived Doppler-boosting factor of, distinguishes CSOs from highly beamed sources such as blazars (Kovalev et al., 2005; Jorstad et al., 2005; Homan et al., 2006; Hovatta et al., 2009; Liodakis et al., 2018; Cheng et al., 2020). Unlike flat-spectrum radio quasars with a core K and Doppler factor of (Pushkarev & Kovalev, 2012), our CSO sample consistently falls below these thresholds (Readhead et al., 1996b; Polatidis & Conway, 2003; An & Baan, 2012), reinforcing their interpretation as young, intrinsically symmetric sources. Because CSOs have moderate brightness temperatures and little Doppler boosting, the observed closely tracks intrinsic emissivity, enabling beaming-independent estimates of jet energy density, size, and age.